OpenStudy (anonymous):

Bridget bought a glass pyramid from a gift store. The pyramid, with a height of 4 cm and a square base of side 6 cm, was packed in a box shaped as a cube of side 7 cm. The empty space in the box was packed with paper tissue. How many cubic centimeters of the box were stuffed with paper tissue?

OpenStudy (anonymous):

volume of pyramid =(1/3)* A*h, A = area of base. Hnece V of pyramid=(1/3)*36*4=48. Volume of cube=7*7*7=343. 343-48=295 so 295cm^3

OpenStudy (anonymous):

Thankyou so much!!